Pearl Furniture Company started construction of a combination office and warehouse building for its own use at an estimated cost of $15,000,000 on January 1, 2020. Pearl expected to complete the building by December 31, 2020. Pearl has the following debt obligations outstanding during the construction period. Construction loan-12% interest, payable semiannually, issued December 31, 2019 $6,000,000 Short-term loan-10% interest, payable monthly, and principal payable at maturity on May 30, 2021 4,200,000 Long-term loan-11% interest, payable on January 1 of each year. Principal payable on January 1, 2024 3,000,000 Assume that Pearl completed the office and warehouse building on December 31, 2020, as planned at a total cost of $15,600,000, and the weighted-average amount of accumulated expenditures was $10,800,000. Compute the avoidable interest on this project. (Use interest rates rounded to 2 decimal places, e.g. 7.58% for computational purposes and round final answers to 0 decimal places, e.g. 5,275.)

Avoidable Interest is ?

Compute the depreciation expense for the year ended December 31, 2021. Pearl elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $900,000. (Round answer to 0 decimal places, e.g. 5,275.)

Depreciation Expense is?

Please help me out! Show steps thanks
Answer #1

Requirement 1

The construction loan is specific loan attributed to construction of building and not avoidable.

The other loans of short term and long term nature and they are general loan. This needs to be verified if they can be avoided.

Computation of weighted average interest rate

Principal

Interest

10% short term loan

$        4,200,000

$ 420,000

11% long term loan

$        3,000,000

$ 330,000

$        7,200,000

$ 750,000

weighted average interest rate = 750,000/7200,000 = 10.42%

Computation of avoidable interest :

Weighted average
Accumulated expenditures

Interest rate

Avoidable Interest

A

B

A X B

6,000,000

12%

$    720,000

10,800,000 - 6000,000 = 4800,000

10.42%

$         500,160

$                                         10,800,000

$            1,220,160

Hence, avoidable interest is $ 1,220,160

Requirement 2

Computation of actual interest

Principal

Interest

12% construction loan

$        6,000,000

$            720,000

10% short term loan

$        4,200,000

$               420,000

11% long term loan

$        3,000,000

$               330,000

$      13,200,000

$               1,470,000

As per accounting standard. If, avoidable interest is lower than actual interest, avoidable interest shall be capitalized to the cost of asset.

Total cost of building = 15,600,000 + 1,220,160 = $ 16,820,160

Depreciation expense = (Cost- salvage value)/Useful life = (16,820,160 – 900,000)/30 = $ 530,672

Hence, required depreciation expense = $ 530,672.